a body is so projected in the air that the horizontal range coveredd b he body is equal to the maximum verticle height attained by the body during the motion. Fing the angle of projection?..........sir plz give the reason and thanks alot in advance

Grade:11

1 Answers

VIKASH YADAV
6 years ago
Max. Vertical height=Horizontal rangeMax. Vertical height is at angle of 45° is u²/2gSo, u²/2g=u²sin2ϴ/g=> sin2ϴ= 1/2=> 2ϴ=sin-1(½)=> 2ϴ=π/6=> ϴ=π/12=15°